OpenStudy (anonymous):

HELP FOR A MEDAL! Find the balance in the account. $4,100 principal earning 4%, compounded monthly, after 10 years

OpenStudy (tkhunny):

Have you considered the single deposit accumulation formula?

OpenStudy (anonymous):

\[4100\left(1+\frac{.04}{12}\right)^{10\times 12}\] and a calculator